# A unified framework for equation discovery and dynamic prediction of hysteretic systems

### v1

@@ 3.2 Solver-based internal variable learning
The second step of the proposed framework focuses on solver-based internal variable learning, where the responses of the formulated state-space equations are simulated through an embedded numerical solver rather than by pointwise regression. Unlike pointwise regression methods that approximate the system’s temporal evolution at individual time steps, the solver-based approach enforces temporal consistency by integrating the governing equations over time, thereby capturing the implicit dependence of the internal hysteretic variable on the system states.
To initialize the model parameters, small-amplitude responses are employed to offer an estimate of the initial values of the system parameters. For Case 1: Hysteresis Discovery, to estimate the initial value of $m,c,k$, the internal variable $z(t)$ in Eq.(3) is omitted, and the dynamic motion equation $m\ddot{x}(t)+c\dot{x}(t)+kx(t)=u(t)$ is used to estimate the initial value $m_{0},c_{0},k_{0},$ while the initial value $\alpha_{0}$ is randomly initialized. For Case 2: Full Equation Discovery, the parameters $m_{0},c_{0},k_{0},\alpha_{0}$ are obtained in the same way as in Case 1: Hysteresis Discovery for estimating the initial value $z_{0}$, while the neural network parameters $\theta_{0}$ are randomly initialized. Based on these settings and Eq.(3), the initial value of internal hysteretic variable $z_{0}$ can be expressed as $z_{0}(t)=[u(t)-m_{0}\ddot{x}(t)-c_{0}\dot{x}(t)-k_{0}x(t)]/\alpha_{0}$. This procedure provides a physically consistent initialization for the learnable hysteretic model, improving convergence and stability of solver-based optimization.
@@
In this step, the governing equations are formulated into a state-space representation, within which a learnable hysteretic model is embedded inside a differentiable ODE solver such as RK4. To make this explicit, the system can be written in the following state-space form:
$\frac{d}{dt}\begin{bmatrix}x\\[4.0pt] \dot{x}\\[4.0pt] z\end{bmatrix}=\begin{bmatrix}\dot{x}\\[4.0pt] f_{\theta}\\[4.0pt] g_{\phi}\end{bmatrix}$, (7)
where $x$ denotes the displacement, $\dot{x}$ is the velocity, and $z$ denotes the hysteretic internal variable; $f_{\theta}$ denotes the dynamic motion equation and $g_{\phi}$ represents the hysteretic link equation. Given the known time stamps $t$, the system responses such as $x(t)$, $\dot{x}(t)$, and $\ddot{x}(t)$ (any of which can be derived or integrated from the others), the external excitation $u(t)$, and the initial conditions $(x(0),\dot{x}(0),z(0))$, these quantities are fed into the learnable hysteretic model within the solver to obtain the predicted responses $x_{\text{pred}}$, $\dot{x}_{\text{pred}}$, and $\ddot{x}_{\text{pred}}$. As described previously, $z$ is defined as a learnable parameter within the model. During each iteration, the errors between predicted values and desired measured data are back-propagated through the solver, to update conditions $z(0)$ and system parameters $\theta$, allowing $z_{\text{pred}}$ to be iteratively refined.
We design a loss function that is flexible to the types of available data. As shown in Figures 1 and 2, the formulation is flexible with respect to the available system responses, such that it can accommodate a subset of displacement, velocity, or acceleration measurements. The corresponding loss functions for the two cases are:
@@
and
[…]
where the selction matrix $\mathbf{S}\in\mathbb{R}^{d\times 3}$ selects specific quantities from data $x_{\text{pred}},\dot{x}_{\text{pred}},\ddot{x}_{\text{pred}}]^{T}\in\mathbb{R}^{3\times N}$, compared with available measurements $\mathbf{y}_{m}\in\mathbb{R}^{d\times N}$; $d=1,2,$ or 3 is the number of measured quantities, and $N$ is the the total number of time steps. After convergence, this process yields the predicted system responses $x_{\text{pred}}$, $\dot{x}_{\text{pred}}$, and $\ddot{x}_{\text{pred}}$, also with predicted internal variables $z_{\text{pred}}$ and $\dot{z}_{\text{pred}}$.
Through this step, the internal hysteretic variable is learned directly from the measurements, addressing one of the central difficulties in hysteresis modeling. Importantly, this process avoids restrictive assumptions, such as prescribing a specific functional form for the hysteretic link equation ($g_{\phi}$) or relying on strongly constrained structural priors, and provides essential data-driven guidance for the subsequent Equation discovery, where SR is employed to recover explicit governing equations.
@@ 3.3 Equation discovery
The final step of the framework is dedicated to extract interpretable governing equations. For the hysteretic link equation ($g_{\phi}$), as illustrated in Figure 2, symbolic regression (SR) is applied to the learned trajectories of $\dot{x}_{\text{pred}}$, $z_{\text{pred}}$, and $\dot{z}_{\text{pred}}$ to search for mathematical expressions $\dot{z}=g_{\phi}(\dot{x},z)$, and this can be written as $\mathrm{SR}\left(\dot{z}_{\text{pred}},\dot{x}_{\text{pred}},z_{\text{pred}}\right)$. For dynamic motion equation ($f_{\theta}$), in Case 1: Hysteresis Discovery case, with known equation structure and the estimated values of $m,c,k,\alpha$, the discovered dynamic motion equation is expressed as $f_{\theta}=(u-c\dot{x}-kx-\alpha z)/m$; In Case 2: Full Equation Discovery, using the same SR method, we can obtain the discovered dynamic motion equation $\ddot{x}=f_{\theta}(x,\dot{x},z,u)$, and this can be written as $\mathrm{SR}\left(\ddot{x}_{\text{pred}},x_{\text{pred}},\dot{x}_{\text{pred}},z_{\text{pred}},u\right)$.
A key advantage of SR lies in its flexibility. Unlike approaches such as SINDy that rely on a predefined functional library, SR enables customisation of the search space by incorporating possible operators (e.g., absolute value and exponential) and explicit control over expression complexity. In particular, SR allows the specification of operator forms without fixing their exact exponents or coefficients – for example, one can include $|\cdot|^{n}$ with $n$ treated as a free constant to be learned. This stands in contrast to SINDy, which requires enumerating many such candidate terms in the predefined library, making it impractical or unfriendly when $n$ is non-integer. Our subsequent experiments will further illustrate this difference. This adaptability mitigates overfitting, promotes interpretability, and facilitates the recovery of compact, physically plausible differential equations that reliably describe the system dynamics.
@@ 4 Experiments
To evaluate the effectiveness and robustness of the proposed framework, we apply the framework to a series of experiments on representative hysteretic systems. These experiments aim to assess both predictive accuracy and generalization ability under varying excitation types, noise conditions and initial conditions. For the noise conditions part, robustness with respect to measurement uncertainty is examined by adding zero-mean Gaussian white noise to measured data, with signal-to-noise ratio (SNR) defined as:
$\text{SNR}(dB)=10\log_{10}\left(\frac{P_{\text{signal}}}{P_{\text{noise}}}\right),$ (9)
@@
The subsequent subsections present three case studies of increasing complexity: (i) the Bouc-Wen hysteretic system benchmark, (ii) a complex structure with nonlinear and fractional terms, and (iii) an experimental SDOF yielding structure. Collectively, these experiments provide a comprehensive evaluation of the unified framework across challenging modeling scenarios.
@@ 4.1 Bouc-Wen hysteretic system benchmark
The Bouc-Wen hysteretic benchmark, developed by Noël and Schoukens [noel2016hysteretic], provides a well-established testbed for evaluating nonlinear system identification methods under hysteretic dynamics. The benchmark builds on the Bouc-Wen model of hysteresis [bouc1967forced, wen1976method], which has been extensively used in structural dynamics and control applications [ismail2009hysteresis, ikhouane2007systems]. The benchmark data are generated through accurate integration of the Bouc-Wen equations using a Newmark scheme, including estimation data obtained from noisy multisine excitations for model identification, as well as separate fixed test datasets for validation. In particular, two fixed test datasets are provided: (i) a sinesweep excitation with amplitude of 40 N sweeping from 20 - 50 Hz and (ii) a random-phase multisine excitation covering the 5 - 150 Hz frequency band [noel2016hysteretic]. These datasets are widely adopted in the system identification community for assessing generalization capabilities of nonlinear models.
@@
The Bouc-Wen hysteretic system can be described by a SDOF nonlinear mass-spring-damper oscillator, where the restoring force includes both elastic and hysteretic components. Following the benchmark description in [noel2016hysteretic], the dynamics are governed by the coupled differential equations:
@@
where $x(t)$ is the displacement response, $u(t)$ is the external excitation, and $z(t)$ is the internal hysteretic variable. The parameters $m$, $c$, and $k$ denote the mass, damping, and linear stiffness, respectively, while $\alpha$, $A$, $\beta$, $\gamma$, and $n$ control the amplitude, shape, and smoothness of the hysteresis loop. For the benchmark configuration, the parameters are set to $m=2.0,c=10.0,k=50000.0,\alpha=1.0,A=50000.0,\beta=800.0,\gamma=-1100.0$ and $n=1.0$.
As the first benchmark case, Figure 3 illustrates the two external excitations employed for training and testing: a sinesweep signal and a multisine signal. In addition, we further consider noisy scenarios by adding Gaussian noise with signal-to-noise ratios of 30 dB, 25 dB, and 20 dB, corresponding to low, medium, and high noise levels, respectively.
Figure 3: Different types of external excitation of benchmark data. The sinesweep signal (left) is used as the training input, while the multisine signal (right) serves as the testing input.
@@
In this benchmark example, we assume to know the structure of the dynamics motion equation, thus it is a problem of Hysteresis Discovery, defined in Figure 1. Based on this set-up, Figure 4 shows the hysteresis loops learned by the proposed framework, in terms of three different representations. It can be observed that, even in the presence of strong noise, the proposed model successfully captures nonlinear memory-dependent hysteretic behaviors.
We define two different types of testing cases. In Testing 1, the external excitation is the same as in Training, but the data are taken from a time interval beyond the Training window. In Testing 2, by contrast, a different external excitation is applied, providing an out-of-distribution scenario relative to Training.
@@
Table 1: Discovered equations results of benchmark data ($t$ is omitted for brevity). The true equations are $2\ddot{x}+10\dot{x}+50000x+z=u$ and $\dot{z}=50000\dot{x}-800|\dot{x}|z+1100\dot{x}|z|$.
Noise
Discovered equations
Noise-free
$1.9998\ddot{x}+10.0231\dot{x}+49893.8617x+1.0110z=u$
$\dot{z}=49633.1496\dot{x}-779.1857|\dot{x}|z+1049.3318\dot{x}|z|$
30dB
$1.9552\ddot{x}+10.0124\dot{x}+49063.4476x+0.8907z=u$
$\dot{z}=60143.4371\dot{x}-811.8583|\dot{x}|z+811.8583\dot{x}|z|$
25dB
$1.8623\ddot{x}+10.2269\dot{x}+47507.9528x+1.0210z=u$
$\dot{z}=50858.0627\dot{x}-734.82275|\dot{x}|z+922.7398\dot{x}|z|$
20dB
$1.7976\ddot{x}+12.8240\dot{x}+46127.6819x+1.2557z=u$
$\dot{z}=43574.2517\dot{x}-727.0708|\dot{x}|z+727.0708\dot{x}|z|$
@@
Noise Responses
Training
NRMSE
Testing 1
NRMSE
Testing 2
NRMSE
@@
To further assess the validity of the discovered models, Figures 5 and 6 compare the displacement and velocity responses obtained by numerically solving the discovered governing equations against the ground-truth responses, respectively. The discovered dynamic equations are numerically solved to obtain the solution trajectories, which are then compared with the benchmark data. This procedure more directly highlights the effectiveness of equation discovery. This benchmark experiment demonstrates that the unified three-step framework is capable of reliably identifying hysteretic dynamics, even under noisy conditions. These findings validate the effectiveness of the proposed methodology for equation discovery and dynamic prediction of nonlinear hysteretic systems.
Figure 5: Displacement results of benchmark data. (a) Training results. (b) Testing 1 (same excitation) results. (c) Testing 2 (different excitation) results.
Figure 6: Velocity results of benchmark data. (a) Training results. (b) Testing 1 (same excitation) results. (c) Testing 2 (different excitation) results.

### v2

@@ 3.2 Solver-based hysteretic variable learning
Step 1 of the proposed framework focuses on solver-based hysteretic variable learning, in which the responses to the are simulated using an embedded numerical solver rather than via pointwise regression. Unlike pointwise regression methods that approximate the system’s temporal evolution at individual time steps, the solver-based approach enforces temporal consistency by integrating the governing equations over time, thereby capturing the implicit dependence of the hysteretic variable on the system states.
Trainable vector variable $\mathbf{z}$. In both cases, the structure of the hysteretic link equation ($g_{\phi}$) is unknown, and the hysteretic variable $\mathbf{z}$ is unobserved. Therefore, we treat $\mathbf{z}$ as a trainable hysteretic vector variable, which is optimized jointly with the model parameters $\theta$. This design is because $g_{\phi}$ has a more complex expression, such as non-smooth operators, which is difficult to represent accurately with a neural network, whereas only the value of the vector $\mathbf{z}$ is required in this step. Therefore, treating $\mathbf{z}$ as a trainable parameter provides a more flexible and accurate formulation.
Initialization. To initialize the model parameters, small-amplitude responses are used to estimate the of $\theta$ and $\mathbf{z}$. Table LABEL:tab:cases shows the initialization of two cases.
For Case 1, $z(t)$ in Eq.(3) is omitted to simplify dynamic motion equation to:
$m\ddot{x}(t)+c\dot{x}(t)+kx(t)=u(t).$ (9)
Then, the initial values of $m,c,k$ are estimated by the least-squares method, while the initial value $\alpha$ is randomly initialized.
For Case 2, the parameters $m,c,k,\alpha$ are obtained in the same way as in Case 1 for estimating the initial value $\mathbf{z}$, and the neural network parameters $\theta$ are randomly initialized.
For the initialization of $\mathbf{z}=[z(t_{1}),z(t_{2}),\ldots,z(t_{N})]^{T}$, where $N$ denotes the total number of sampled time points, based on Eq.(3), we can express as:
$z(t)=[u(t)-m\ddot{x}(t)-c\dot{x}(t)-kx(t)]/\alpha$ (10)
and get the initial $\mathbf{z}$. This procedure provides a proper physically consistent initialization for the learnable hysteretic model, improving convergence and stability of solver-based optimization.
@@
Embedded Solver with trainable $\theta$ and $\mathbf{z}$. In this step, the governing equations are reformulated into a state-space representation, within which learnable hysteretic model is embedded inside a differentiable ODE solver such as RK4.
Given the known time stamps $t_{i}$ ($i=1,2,\ldots,N$), the time step $\Delta t$, the corresponding input $u_{i}$, and the initial state $x(t_{1}),\dot{x}(t_{1})]^{T}$, the state trajectory can be computed sequentially using the ODE solver as:
@@
From above solver step, it can be seen that each element in the vector $\mathbf{z}=[z(t_{1}),z(t_{2}),\ldots,z(t_{N})]^{T}$, together with $\theta$, is used within the solver to generate the predicted responses $\mathbf{x}_{\text{pred}}$, $\mathbf{\dot{x}}_{\text{pred}}$, and $\mathbf{\ddot{x}}_{\text{pred}}$, and is continuously updated during the training iterations. In this way, the solver is embedded into the subsequent training process.
Training. Based on the above embedded solver formulation, the training inputs are given by $\{[x(t_{1}),\dot{x}(t_{1})]^{T},\,f_{\theta},\,\mathbf{u},\,\mathbf{z}\}$ ($x(t_{1}),\dot{x}(t_{1})]^{T}$ is the initial condition), and the solver produces the predicted responses $\mathbf{x}_{\text{pred}}$, $\mathbf{\dot{x}}_{\text{pred}}$, and $\mathbf{\ddot{x}}_{\text{pred}}$. In both cases, training is performed by backpropagation to update the parameters $\theta$ and $\mathbf{z}$.
We design a loss function that is flexible to the types of available data. Given the known time stamps $t_{i},i=1,2,\ldots,N$, the corresponding system responses such as $x(t_{i})$, $\dot{x}(t_{i})$, and $\ddot{x}(t_{i})$ (which may include a subset of them depending on sensing). As shown in Figures 1 and 2, the formulation is flexible with respect to the available system responses. The corresponding loss function is:
$\begin{array}[]{@{}l@{}}\mathcal{L}\text{oss}_{\theta,\mathbf{z}}\bigl(\mathbf{x}_{\text{pred}},\mathbf{\dot{x}}_{\text{pred}},\mathbf{\ddot{x}}_{\text{pred}}\bigr)=\left\lVert\,\mathbf{S}\,\begin{bmatrix}\mathbf{x}_{\mathrm{pred}}\\ \mathbf{\dot{x}}_{\mathrm{pred}}\\ \mathbf{\ddot{x}}_{\mathrm{pred}}\end{bmatrix}-\mathbf{Y}\right\rVert_{2}^{2}\end{array},$ (12)
where the selection matrix $\mathbf{S}$ specifies which quantities are observed and $\mathbf{Y}$ denotes the corresponding available measurements (the same definitions as Figure 1).
In practice, when some response components are not directly measured, they can be reconstructed from the available signals through smoothing, numerical differentiation, or numerical integration. In our examples, this means that the full set of responses is numerically completed, so that $\mathbf{S}$ is an identity matrix. More generally, in our framework, $\mathbf{S}$ can be specified flexibly according to the available observations, because $x(t_{i})$, $\dot{x}(t_{i})$, and $\ddot{x}(t_{i})$ can all be incorporated into the optimization of $\theta$ and $\mathbf{z}$. This is one aspect of the flexibility of the proposed framework.
Once training is completed, the predicted system responses $\mathbf{x}_{\text{pred}}$, $\mathbf{\dot{x}}_{\text{pred}}$, and $\mathbf{\ddot{x}}_{\text{pred}}$, together with the predicted hysteretic variable $\mathbf{z}_{\text{pred}}$, are obtained. The corresponding $\mathbf{\dot{z}}_{\text{pred}}$ is then computed via finite differences, denoted $\mathrm{Diff}(\mathbf{z}_{\text{pred}})$.
Through this step, the hysteretic variable is learned directly from measurements, addressing one of the central difficulties in hysteresis modeling. Importantly, this process avoids restrictive assumptions, such as prescribing a specific functional form for the hysteretic link equation ($g_{\phi}$) or relying on strongly constrained structural priors, and provides essential data-driven guidance for the subsequent Step 2: Equation discovery, where SR is employed to recover explicit governing equations.
@@ 3.3 Equation discovery
Step 2 of the framework is dedicated to extracting interpretable governing equations. For the hysteretic link equation ($g_{\phi}$), as illustrated in Figure 2, symbolic regression (SR) is applied to the learned trajectories of $\dot{x}_{\text{pred}}$, $z_{\text{pred}}$, and $\dot{z}_{\text{pred}}$ to search for mathematical expressions $\dot{z}=g_{\phi}(\dot{x},z)$, and this can be written as $\mathrm{SR}\left(\dot{z}_{\text{pred}},\dot{x}_{\text{pred}},z_{\text{pred}}\right)$. For dynamic motion equation ($f_{\theta}$), in Case 1: Hysteresis Discovery, with known equation structure and the estimated values of $m,c,k,\alpha$, the discovered dynamic motion equation is expressed as $f_{\theta}=(u-c\dot{x}-kx-\alpha z)/m$; in Case 2: Full Equation Discovery, using the same SR method like $g_{\phi}$, we can obtain the discovered dynamic motion equation $\ddot{x}=f_{\theta}(x,\dot{x},z,u)$, and this can be written as $\mathrm{SR}\left(\ddot{x}_{\text{pred}},x_{\text{pred}},\dot{x}_{\text{pred}},z_{\text{pred}},u\right)$.
A key advantage of SR lies in its flexibility. Unlike approaches such as SINDy that rely on a predefined functional library, SR enables customization of the search space by incorporating possible operators (e.g., absolute value and exponential) and explicit control over expression complexity. In particular, SR allows the specification of operator forms without fixing their exact exponents or coefficients – for example, one can include $|\cdot|^{n}$ with $n$ treated as a free constant to be learned. This stands in contrast to SINDy, which requires enumerating many such candidate terms in the predefined library, making it impractical or unfriendly when $n$ is non-integer. Our subsequent experiments will further illustrate this difference. This adaptability mitigates overfitting, promotes interpretability, and facilitates the recovery of compact, physically plausible differential equations that reliably describe the system dynamics.
@@ 4 Experiments
To evaluate the effectiveness and robustness of the proposed framework, we apply the framework to a series of experiments on representative hysteretic systems. These experiments aim to assess both predictive accuracy and generalization ability under varying excitation types, noise conditions and initial conditions. For the noise conditions part, robustness with respect to measurement uncertainty is examined by adding zero-mean Gaussian white noise to the measured data, with signal-to-noise ratio (SNR) defined as:
$\text{SNR}(\text{dB})=10\log_{10}\left(\frac{P_{\text{signal}}}{P_{\text{noise}}}\right),$ (13)
@@
The subsequent subsections present case studies of increasing complexity: (i) the Bouc-Wen hysteretic system benchmark, (ii) a complex structure with nonlinear and fractional terms, (iii) an experimental single degree-of-freedom (SDOF) yielding structure, and . Collectively, these experiments provide a comprehensive evaluation of the proposed framework.
@@ 4.1 Bouc-Wen hysteretic system benchmark
The Bouc-Wen hysteretic benchmark, developed by Noël and Schoukens [32], provides a well-established testbed for evaluating nonlinear system identification methods under hysteretic dynamics. The benchmark builds on the Bouc-Wen model of hysteresis [5, 43], which has been extensively used in structural dynamics and control applications [19, 17]. The benchmark data are generated through accurate integration of the Bouc-Wen equations using a Newmark scheme, including estimation data obtained from noisy multisine excitations for model identification, as well as separate fixed test datasets for validation. In particular, two fixed test datasets are provided: (i) a sinesweep excitation with amplitude of 40 N sweeping from 20 - 50 Hz and (ii) a random-phase multisine excitation covering the 5 - 150 Hz frequency band [32]. These datasets are widely adopted in the system identification community for assessing generalization capabilities of nonlinear models.
@@
The Bouc-Wen hysteretic system can be described by a SDOF nonlinear mass-spring-damper oscillator, where the restoring force includes both elastic and hysteretic components. Following the benchmark description in [32], the dynamics are governed by the coupled differential equations:
@@
where $x(t)$ is the displacement response, $u(t)$ is the external excitation, and $z(t)$ is the hysteretic variable. The parameters $m$, $c$, and $k$ denote the mass, damping, and linear stiffness, respectively, while $\alpha$, $A$, $\beta$, $\gamma$, and $n$ control the amplitude, shape, and smoothness of the hysteresis loop. For the benchmark configuration, the parameters are set to $m=2.0,c=10.0,k=50000.0,\alpha=1.0,A=50000.0,\beta=800.0,\gamma=-1100.0$ and $n=1.0$.
As the first benchmark case, In addition, we further consider noisy scenarios by adding Gaussian noise with signal-to-noise ratios of 30 dB, 20 dB, and dB, corresponding to low, medium, and high noise levels, respectively.
Figure 3: Different types of external excitation of benchmark data. The sinesweep signal (left) is used as the training input (Training and Testing 1), while the multisine signal (right) serves as the testing input (Testing 2).
@@
Here, we assume to know the structure of the dynamics motion equation, thus it is a problem of Hysteresis Discovery. Figure 4 shows the hysteresis loops learned by the proposed framework, in terms of three different representations. It can be observed that, even in the presence of strong noise, the proposed model successfully captures nonlinear memory-dependent hysteretic behaviors.
We define two different types of testing cases. In Testing 1, the external excitation is the same as in Training, but the data are taken from a time span beyond the training window. In Testing 2, an external excitation is applied, providing an out-of-distribution scenario relative to training.
@@
Table 1: Discovered equations results of benchmark data ($t$ is omitted for brevity).
Noise Discovered equations
Noise-free $1.9998\ddot{x}+10.0231\dot{x}+49893.8617x+1.0110z=u$ $\dot{z}=49633.1496\dot{x}-779.1857|\dot{x}|z+1049.3318\dot{x}|z|$
30dB $1.8623\ddot{x}+10.2269\dot{x}+47507.9528x+1.0210z=u$ $\dot{z}=50858.0627\dot{x}-734.82275|\dot{x}|z+922.7398\dot{x}|z|$
20dB $1.7976\ddot{x}+12.8240\dot{x}+46127.6819x+1.2557z=u$ $\dot{z}=43574.2517\dot{x}-727.0708|\dot{x}|z+727.0708\dot{x}|z|$
@@
Noise Responses Training NRMSE Testing 1 NRMSE Testing 2 NRMSE
@@
To further evaluate the validity of the discovered models, This procedure more directly highlights the effectiveness of equation discovery. This benchmark experiment demonstrates that the unified framework is capable of reliably identifying hysteretic dynamics, even under noisy conditions. These findings validate the effectiveness of the proposed methodology for equation discovery and dynamic prediction of nonlinear hysteretic systems.
Figure 5: Displacement results of benchmark data. (a) Training results. (b) Testing 1 (same excitation) results. (c) Testing 2 ( excitation) results.
Figure 6: Velocity results of benchmark data. (a) Training results. (b) Testing 1 (same excitation) results. (c) Testing 2 ( excitation) results.
To further validate the proposed framework, we also conduct some additional benchmark experiments beyond the baseline setting: Cross-excitation validation, Extrapolation test on unseen amplified excitation, Sensitivity analyses, Noisy wideband random excitation training and Without hysteresis. The detailed context can be seen in C.